# Convolutional Monge Mapping between EEG Datasets to Support Independent Component Labeling

### v1

@@ Abstract
EEG recordings contain rich information about neural activity but are subject to artifacts, noise, and superficial differences due to sensors, amplifiers, and filtering. Independent component analysis and automatic labeling of independent components (ICs) enable artifact removal in EEG pipelines. Convolutional Monge Mapping Normalization (CMMN) is a recent tool used to achieve spectral conformity of EEG signals, which was shown to improve deep neural network approaches for sleep staging. Here we propose a novel extension of the CMMN method with two alternative approaches to computing the source reference spectrum the target signals are mapped to: (1) channel-averaged and $l_{1}$-normalized barycenter, and (2) a subject-to-subject mapping that finds the source subject with the closest spectrum to the target subject. Notably, our extension yields space-time separable filters that can be used to map between datasets with different numbers of EEG channels. We apply these filters in an IC classification task, and show significant improvement in recognizing brain versus non-brain ICs.
Clinical relevance—EEG recordings are used in the diagnosis and monitoring of multiple neuropathologies, including epilepsy and psychosis. While EEG analysis can benefit from automating artifact removal through independent component analysis and labeling, differences in recording equipment and context (the presence of noise from electrical wiring and other devices) may impact the performance of machine learning models, but these differences can be minimized by appropriate spectral normalization through filtering.
I INTRODUCTION
In machine learning, out-of-distribution data (data from different domains) can deteriorate performance of models, especially when trained on limited data. The model may be sensitive to differences in the data distribution that are not actually relevant to the task. In the context of biomedical data, domains may correspond to different equipment, recording settings, or the population of individuals. In the context of EEG, differences in set-up (electrodes, amplifiers, analog and digital filters, power line noise) can create stark changes in the spectral content of recordings. One method to address this problem is the Convolutional Monge Mapping Normalization (CMMN) approach for spectral normalization based on optimal transport [1], which optimizes a linear filter for each subject and channel such that after filtering all subjects have the same spectrum for that channel, which is the barycenter of the channel’s spectra for all training subjects. Originally, CMMN was applied to sleep staging with deep neural networks.
@@ I INTRODUCTION
In this work, we extend it to the case of automatically labeling independent components, such as the popular ICLabel classifier [2], which can support automatic EEG artifact removal and denoising. In comparison to the ICLabel classifier, which uses a neural network model with scalp map and spectral features (including the autocorrelation sequence), and was trained on an extensive dataset [3], we consider custom classifiers trained on a much smaller dataset collected in the US [4] using only time-series features extracted from each independent component [5]—that is, without any spatial information. We then test on an another independent dataset [6], collected in Europe. These two datasets were used in a prior work on IC classification [7], and are now both available on OpenNeuro [8, 9]. Because these datasets have expert-labeled ICs, we avoid filtering each channel with a different filter, which would change each IC differently; instead, we find a single CMMN filter for each subject that is applied to all channels. This channel-averaged CMMN filter is based on matching the channel-averaged spectrum of each test subject to either the barycenter of the channel-averaged spectra from the training set or the nearest spectrum in the training set in terms of the optimal transport distance. With a shared single temporal filter across channels, the filtering can be applied before or after the spatial filtering achieved by ICA’s demixing matrix. That is, the joint filtering is separable.
Crucially, the use of the channel-averaging enables the application of CMMN to target domains with a differing number of channels. Here, the source (train) and target (test) datasets have 134–235 channels and 64 channels, respectively.
This extension of the CMMN methodology achieves domain adaptation between two datasets with the pre-trained classifiers such that their performance (the subject averaged F1 score for the brain class is 0.91) exceeds the performance of the ICLabel (0.88).
II METHODOLOGY
We first detail the background of convolutional Monge mapping normalization [1] that was not wholly contained in that work. We then detail the proposed additions.
II-A Background
The CMMN methodology enables domain adaptation between different neural signals [1] by mapping the spectra of the target subjects to the barycenter of the source subjects via a linear filter that achieves the optimal transport (under an $\ell_{2}$-metric) under zero-mean stationary Gaussian discrete-time processes. This follows from the fact that the Wasserstein-2 distance between Gaussian distributions has a closed form [10, 11] known as the Bures-Wasserstein distance or Fréchet distance, expressed in terms of the means $\bm{m}^{\text{S}},\bm{m}^{\text{T}}$ and covariance matrices $\bm{\Sigma}^{\text{S}},\bm{\Sigma}^{\text{T}}$, as
where the second term in the square root is the squared Bures distance. Under the stationarity assumption, a discrete-time zero-mean Gaussian process is completely described by its auto-covariance matrix, which is a symmetric and Toeplitz matrix formed from the auto-correlation sequence $r[k]=\mathbb{E}[x[n]x[n+k]]$. Assuming $r[\tau]$ decays for increasing lags and taking a sufficiently large lag (or applying a windowing function to the auto-correlation sequence as in the Blackman–Tukey method), and optionally padding with $Q$ zeros, the auto-correlation sequence can be formed into an $N\times N$ circulant matrix (necessarily Toeplitz and still symmetric) where
$\bm{r}=[\ r[0],r[1],\ldots,r[K-1],\underbrace{0,\ldots,0}_{Q},r[-K+1],\ldots,r[-1]\ ]$ (2)
is the first column. Let $\bm{\Sigma}_{\bm{r}}$ denote the resulting $N\times N$ circulant matrix, with N=2K1+QN=2K-1+Q. Then, the discrete Fourier transform (DFT) matrix $\bm{F}=[e^{-\mathrm{i}\frac{2\pi kn}{N}}]_{k=1,n=1}^{N,N}$ diagonalizes it, $\bm{F}\bm{\Sigma}_{\bm{r}}\bm{F}^{H}=\mathrm{diag}(\bm{\lambda}_{\bm{r}})$, where $\bm{F}^{H}$ denotes the Hermitian transpose of $\bm{F}$ and $\mathrm{i}=\sqrt{-1}$. The resulting diagonal
$\bm{\lambda}_{\bm{r}}=\mathrm{diag}(\bm{F}\bm{\Sigma}_{\bm{r}}\bm{F}^{H})=\bm{F}\bm{r}$ (3)
corresponds to an statistical estimate of the power spectral density (PSD). In contrast, a finite-window periodogram estimate of the PSD is
$\bm{\lambda}$ $=\mathbb{E}[(\bm{F}\bm{x})^{\odot 2}]=\mathbb{E}[\mathrm{diag}(\bm{F}\bm{x}\bm{x}^{H}\bm{F}^{H})]$ (4)
$=\mathrm{diag}(\bm{F}\mathbb{E}[\bm{x}\bm{x}^{H}]\bm{F}^{H})=\mathrm{diag}(\bm{F}\bm{\Sigma}\bm{F}^{H}),$
where $\bm{x}$ is a window of length $N$ from the process and $\odot$ denotes elementwise power. Note that clearly $\bm{\lambda}$ and $\bm{\Sigma}=\mathbb{E}[\bm{x}\bm{x}^{H}]$ are non-negative and positive semidefinite, respectively. $\bm{\Sigma}$ will only be circulant under the assumption of periodic discrete-time signals such that $\bm{\Sigma}=\bm{\Sigma}_{\bm{r}}$ and $\bm{\lambda}=\bm{\lambda}_{\bm{r}}$; otherwise, $\bm{F}\bm{\Sigma}\bm{F}^{H}$ will not be strictly diagonal. The Wiener-Khinchin theorem, equating the Fourier transform of the autocorrelation with the power spectral density, only holds for infinitely auto-correlation sequences. Truncating the auto-correlation sequence, without proper choice of a windowing function, can lead to a circulant matrix $\bm{\Sigma}_{\bm{r}}$ which is not positive semidefinite, yielding negative values in the spectrum. The truncation corresponds to a rectangular window whose Fourier transform is a sinc function, which takes positive and negative values. Multiplication in time-domain causes convolution in the frequency domain and the result may be negative. Using a windowing function whose Fourier transform is strictly positive, such as the Barlett (triangular), Parzen, or Gaussian, will ensure a non-negative spectrum. In fact, the finite-window periodogram estimate (4) uses a Barlett window.
The squared Wasserstein-2 distance between two discrete-time, zero-mean stationary Gaussian random processes with auto-covariance $\bm{\Sigma}^{\text{S}}$ and $\bm{\Sigma}^{\text{T}}$, which are circulant, is
$W_{2}(\mathcal{N}(\bm{0},\bm{\Sigma}^{\text{S}}),\mathcal{N}(\bm{0},\bm{\Sigma}^{\text{T}}))=\lVert\bm{\lambda}^{\text{S}\odot\frac{1}{2}}-\bm{\lambda}^{\text{T}\odot\frac{1}{2}}\rVert_{2},$ (5)
using the fact that if two matrices share the same basis, then the Bures distance is the Euclidean distance between the square root of their spectra. This can be seen directly by substituting $\bm{\Sigma}^{\text{S}}=\bm{F}^{H}\mathrm{diag}(\bm{\lambda}^{\text{S}})\bm{F}$ and $\bm{\Sigma}^{\text{T}}=\bm{F}^{H}\mathrm{diag}(\bm{\lambda}^{\text{T}})\bm{F}$ into the squared Bures distance and using $\bm{F}^{H}\bm{F}=\bm{I}$:
$\mathrm{tr}$ $(\bm{\Sigma}^{\text{S}}+\bm{\Sigma}^{\text{T}}-(\bm{\Sigma}^{\text{S}\frac{1}{2}}\bm{\Sigma}^{\text{T}}\bm{\Sigma}^{\text{S}\frac{1}{2}})^{\frac{1}{2}})$ (6)
$=\mathrm{tr}(\bm{F}^{H}\mathrm{diag}(\bm{\lambda}^{\text{S}})\bm{F}+\bm{F}^{H}\mathrm{diag}(\bm{\lambda}^{\text{T}})\bm{F}$
$\ \quad-(\bm{F}^{H}\mathrm{diag}(\bm{\lambda}^{\text{S}\odot\frac{1}{2}})\mathrm{diag}(\bm{\lambda}^{\text{T}})\mathrm{diag}(\bm{\lambda}^{\text{S}\odot\frac{1}{2}})\bm{F})^{\frac{1}{2}})$
$=\mathrm{tr}(\mathrm{diag}(\bm{\lambda}^{\text{S}})+\mathrm{diag}(\bm{\lambda}^{\text{T}})-(\bm{F}^{H}\mathrm{diag}(\bm{\lambda}^{\text{S}}\odot\bm{\lambda}^{\text{T}})\bm{F})^{\frac{1}{2}})$
$=\mathrm{tr}(\mathrm{diag}(\bm{\lambda}^{\text{S}})+\mathrm{diag}(\bm{\lambda}^{\text{T}})-\bm{F}^{H}\mathrm{diag}(\bm{\lambda}^{\text{S}\frac{1}{2}}\odot\bm{\lambda}^{\text{T}\frac{1}{2}})\bm{F})$
$=\sum_{n=0}^{N-1}\bm{\lambda}^{\text{S}}[n]+\bm{\lambda}^{\text{T}}[n]-\sqrt{\bm{\lambda}^{\text{S}}[n]\bm{\lambda}^{\text{T}}[n]}$
$=\lVert\bm{\lambda}^{\text{S}\odot\frac{1}{2}}-\bm{\lambda}^{\text{T}\odot\frac{1}{2}}\rVert_{2}^{2},$
where $\odot$ also denotes the elementwise (Hadamard) product.
In the case, that all the covariance matrices are circulant and positive semidefinite, it is straightforward to show that $\bm{\Sigma}_{\text{S}}=\bm{F}^{H}\mathrm{diag}(\bar{\bm{\lambda}}_{\text{S}})\bm{F}$, where $\bar{\bm{\lambda}}_{\text{S}}=(\frac{1}{I}\sum_{i}\bm{\lambda}_{i}^{\text{S}\odot\frac{1}{2}})^{\odot 2}$. Using (5),
$\sum_{i=1}^{I}\frac{1}{I}W_{2}^{2}(\mathcal{N}(\bm{0},\bm{\Sigma}),\mathcal{N}(\bm{0},\bm{\Sigma}_{i}^{\text{S}})=\sum_{i=1}^{I}\frac{1}{I}\lVert\bm{\lambda}^{\odot\frac{1}{2}}-\bm{\lambda}_{i}^{\text{S}\odot\frac{1}{2}}\rVert_{2}^{2}.$ (8)
Taking the derivative with respect to $\bm{\lambda}^{\odot\frac{1}{2}}$ yields $\bm{\lambda}^{\odot\frac{1}{2}}=\frac{1}{I}\sum_{i}\bm{\lambda}_{i}^{\text{S}\odot\frac{1}{2}}$.
$\bm{\lambda}^{y}$ $=\mathbb{E}[|\bm{F}\bm{y}|^{\odot 2}]=\mathbb{E}[|(\bm{F}\bm{x})\odot\bm{H}|^{\odot 2}]$ (9)
$=\mathbb{E}[|\bm{F}\bm{x}|^{\odot 2}\odot|\bm{H}|^{\odot 2}]=|\bm{H}|^{\odot 2}\odot\mathbb{E}[|\bm{F}\bm{x}|^{\odot 2}]$ (10)
$=|\bm{H}|^{\odot 2}\odot\bm{\lambda}^{\text{T}},$ (11)
Clearly, the distance is zero for any filter $h[n]$ such that
$|\bm{H}|=\bm{\lambda}^{\text{S}\odot\frac{1}{2}}\odot\bm{\lambda}^{\text{T}\odot-\frac{1}{2}}\in\mathbb{R}_{\geq 0}^{N},$ (13)
i.e., where the magnitude of the frequency response is the square-root of the power spectral densities $|H[n]|=\sqrt{\frac{{\lambda}^{\text{S}}[n]}{\lambda^{\text{T}}[n]}},n\in\{0,\ldots,N-1\}$.
II-B Proposed Methodology
@@
We extend this methodology in two key ways: we propose to use channel-averaged PSDs yielding a single filter per subject and a subject-to-subject mapping scheme. A key benefit of using a single filter per subject is that it enables adapting between EEG datasets with a variable number of channels. We also propose to use $\ell_{1}$-normalization of spectra for invariance to signal scale, which may occur due to impedance differences due to electrodes or scalp contact or differences in the electronic systems (amplifiers and filtering) or digital filtering.
The $\ell_{1}$-normalized PSD is a probability mass function (PMF)—non-negative and sums to 1. Let $\tilde{\bm{\lambda}}^{\text{S}}=\frac{\bm{\lambda}^{\text{S}}}{\lVert\bm{\lambda}^{\text{S}}\rVert_{1}}$, this corresponds to the circulant matrix $\bm{\rho}^{\text{S}}=\frac{\bm{\Sigma}^{\text{S}}}{\mathrm{tr}(\bm{\Sigma}^{\text{S}})}$, and likewise for the target. Then, the Wasserstein-2 distance between two discrete-time zero-mean variance-normalized stationary Gaussian random processes is
$W_{2}($ $\mathcal{N}(\bm{0},\bm{\rho}^{\text{S}}),\mathcal{N}(\bm{0},\bm{\rho}^{\text{T}}))=\lVert\tilde{\bm{\lambda}}^{\text{S}\odot\frac{1}{2}}-\tilde{\bm{\lambda}}^{\text{T}\odot\frac{1}{2}}\rVert_{2}$ (14)
$=\sqrt{2-2\langle\tilde{\bm{\lambda}}^{\text{S}\odot\frac{1}{2}},\tilde{\bm{\lambda}}^{\text{T}\odot\frac{1}{2}}\rangle}$ (15)
$=\sqrt{2}\sqrt{1-\langle\tilde{\bm{\lambda}}^{\text{S}\odot\frac{1}{2}},\tilde{\bm{\lambda}}^{\text{T}\odot\frac{1}{2}}\rangle}=\sqrt{2}d_{\text{He}}(\tilde{\bm{\lambda}}^{\text{S}},\tilde{\bm{\lambda}}^{\text{T}}),$ (16)
where $d_{\text{He}}$ denotes the Hellinger distance, which is applicable since the $\ell_{1}$-normalized PSDs reside on the probability simplex $\Delta^{N}=\{\tilde{\bm{\lambda}}\in\mathbb{R}_{\geq 0}^{N}:\sum_{n=0}^{N-1}\tilde{\lambda}[n]=1\}$. The use of Hellinger distance compared to Bures-Wasserstein distance means that signals are compared in terms of their spectral shape, without regard to their variance.
Given this background, we now proceed to consider access to data (multichannel EEG) from $I$ subjects in the source domain and one or more target subjects. We apply CMMN to transform the target EEGs to the source domain and apply classifiers trained on the source domain to the transformed target data. Let $x_{c}[n]$ denote the $c$-th channel of the target signal, $c\in\{1,\ldots,C^{\text{T}}\}$. The CMMN transformation is a linear filtering of $x_{c}[n]$:
$y_{c}[n]=(h*x_{c})[n],\quad c\in\{1,\ldots,C^{\text{T}}\},$ (17)
@@
where $h[n]$ is the normalizing CMMN filter, and $y_{c}[n]$ is the transformed signal that will have a source-like spectrum (the channel-averaged PSDs will match). The estimation of the normalizing CMMN filter is a three-step process that uses a reference spectrum (computed in the second step) that is either a Barycenter from the source or the normalized PSD of a source subject via the subject-to-subject (Subj-to-subj) assignment:
  1. 1)
    Power spectral density calculation
  2. 2a)
    Barycenter calculation
  3. 2b)
    Minimization of the Hellinger distance between target signals and source signals for the Subj-to-subj mapping scheme
  4. 3)
    Calculation of the normalizing filters
@@
Step 1 PSD calculation. As in [1], we use the Welch periodogram method [14], which takes averages of the squared FFT from possibly overlapping windows, to calculate PSDs for each subject, as implemented by the SciPy library [15]. For real-valued signals, the spectra will be conjugate symmetric for positive and negative frequencies. In this case, $\bm{p}\in\mathbb{R}_{\geq 0}^{P}$ denotes the PSD for $P$ non-negative frequencies (for zero-mean signals the DC component should be zero but is retained). For $K$ equal-sized windows of length $M$, $\{\bm{x}_{k}\}_{k=1}^{K}=\mathcal{X}\subset\mathbb{R}^{M}$ with windowing function $\bm{w}$, the power spectral density estimate is
$\bm{p}=\frac{1}{|\mathcal{X}|}\sum_{\bm{x}\in\mathcal{X}}\lvert\mathrm{RFFT}(\bm{w}\odot\bm{x})\rvert^{\odot 2}\in\mathbb{R}_{\geq 0}^{P},$ (18)
where $\mathrm{RFFT}$ denotes the FFT operation for real-valued signals and returns the density estimate only for non-negative frequencies, $P=\mathtt{nfft}/2+1$ is the number of non-negative frequencies, and $N=\mathtt{nfft}$ is the length of the windows after zero-padding, assumed to be even.
For a subject with $C$ channels, the matrix of PSDs is $\bm{P}=[\bm{p}_{1},\ldots,\bm{p}_{C}]^{\top}\in\mathbb{R}^{C\times P}$. Unlike the original CMMN, which used different filters for each channel, we propose to compute a channel-averaged PSD as $\bar{\bm{p}}=\frac{1}{C_{i}}\sum_{c=1}^{C}\bm{p}_{c}$.
@@
Step 2a Barycenter calculation. We compute a barycenter for the source subjects only, this serves as the template for mapping the target subjects. The barycenter for $I$ source subjects $\bar{\bm{p}}_{1}^{\text{S}},\ldots,\bar{\bm{p}}_{I}^{\text{S}}$ channel-averaged PSDs is defined as
${\bar{\bm{p}}}_{\text{S}}=\frac{1}{I}\sum_{i=1}^{I}\bar{\bm{p}}^{\text{S}}_{i}.$ (19)
Additionally, for robustness against differing scales we propose an option for calculating a normalized barycenter for the source subjects, where each source subject PSD is normalized, as
${\tilde{\bm{p}}}_{\text{S}}=\frac{1}{I}\sum_{i=1}^{I}\frac{\bar{\bm{p}}^{\text{S}}_{i}}{\|\bar{\bm{p}}^{\text{S}}_{i}\|_{1}}=\frac{1}{I}\sum_{i=1}^{I}\tilde{\bm{p}}^{\text{S}}_{i},\text{ where }\tilde{\bm{p}}^{\text{S}}_{i}=\frac{\bar{\bm{p}}^{\text{S}}_{i}}{\|\bar{\bm{p}}^{\text{S}}_{i}\|_{1}}.$ (20)
@@
This ensures that each subject contributes equally to the average. Note that as the units of PSD are squared, outliers without normalization can influence the average. However, we avoided channel-wise normalization because that could change the contribution of individual channels, but this option could also be explored.
Step 2b Subject-to-subject mapping scheme for post-hoc alignment. Our proposed alternative to the barycenter mapping scheme associates a given target subject to the closest source subject with respect to the Wasserstein distance between variance-normalized zero-mean stationary Gaussian processes, which as stated is proportional to the Hellinger distance. Let $\tilde{\bm{p}}^{\text{T}}=\frac{\bar{\bm{p}}^{\text{T}}}{\|\bar{\bm{p}}^{\text{T}}\|_{1}}$ denote the $\ell_{1}$-normalized channel-averaged PSD for a particular target subject. The channel-averaged PSD of best matched source subject $\bar{\bm{p}}_{*}^{\text{S}}$ is
$\bar{\bm{p}}_{*}^{\text{S}}$ $=\bar{\bm{p}}_{i^{*}}^{\text{S}},\quad i^{*}=\operatorname*{arg\,min}_{i\in\{1,\ldots,I\}}d_{\text{He}}(\tilde{\bm{p}}^{\text{S}}_{i},\tilde{\bm{p}}^{\text{T}}),$ (21)
where the Hellinger distance between two normalized PSDs $\tilde{\bm{p}}^{\text{S}},\tilde{\bm{p}}^{\text{T}}\in\Delta^{P}\subset[0,1]^{P}$, where $\Delta^{P}=\{\tilde{\bm{p}}\in\mathbb{R}_{\geq 0}^{P}:\sum_{n=0}^{P-1}\tilde{p}[n]=1\}$ denotes the probability simplex, is
@@
$d_{\text{He}}(\tilde{\bm{p}}^{\text{S}},\tilde{\bm{p}}^{\text{T}})$ $=\frac{1}{\sqrt{2}}\lVert\tilde{\bm{p}}^{\text{S}\odot\frac{1}{2}}-\tilde{\bm{p}}^{\text{T}\odot\frac{1}{2}}\rVert_{2}$ (22)
$=\frac{1}{\sqrt{2}}\sqrt{\sum_{n=0}^{P-1}\left(\sqrt{\tilde{p}^{\text{S}}[n]}-\sqrt{\tilde{p}^{\text{T}}[n]}\right)^{2}}.$ (23)
@@
Step 3 Normalizing filter. We define a real-valued, zero-phase linear filter with impulse response $h[n]$ and frequency response $\bm{H}$. This filter solves the optimal transport problem for both the barycenter and subject-to-subject formulations, and its frequency response is given by the square root of the ratio of the channel-averaged PSDs [16]:
$\bm{h}$ $=\mathrm{IRFFT}_{M}(\bm{H}),$ (24)
$\bm{H}$ $=\bar{\bm{p}}^{\text{T}\odot-\frac{1}{2}}\odot\bar{\bm{p}}^{\text{S}\odot\frac{1}{2}},$ (25)
@@
$H[n]$ $=\frac{\sqrt{\bar{p}^{\text{S}}[n]}}{\sqrt{\bar{p}^{\text{T}}[n]}},n\in\{0,\ldots,P-1\},$ (26)

### v2

# Learning from Time-Series for Health

@@ Abstract
EEG recordings contain rich information about neural activity and are used in the diagnosis and monitoring of multiple neuropathologies, including epilepsy and psychosis, but are subject to artifacts and noise. While EEG analysis can benefit from automating artifact removal through independent component analysis and automatic labeling of independent components (ICs), differences in recording equipment and context (the presence of noise from electrical wiring and other devices) may impact the performance of IC classifiers. Here we investigate how these differences can be minimized by appropriate spectral normalization through filtering using Convolutional Monge Mapping Normalization (CMMN), which was previously shown to improve deep neural network approaches for sleep staging. We propose a novel extension of the CMMN method with two alternative approaches to computing the source reference spectrum the target signals are mapped to: (1) channel-averaged and $l_{1}$-normalized barycenter, and (2) a subject-to-subject mapping that finds the source subject with the closest spectrum to the target subject. Notably, our extension yields space-time separable filters that can be used to map between datasets with different numbers of EEG channels. We apply these filters in an IC classification task, and show significant improvement in recognizing brain versus non-brain ICs.
@@
In the context of EEG, differences in set-up (electrodes, amplifiers, analog and digital filters, power line noise) can create stark changes in the spectral content of recordings, which can deteriorate the performance of models, especially when trained on limited data. A solution to this problem is the Convolutional Monge Mapping Normalization (CMMN) approach, a method for spectral normalization based on optimal transport [7]. The CMMN method optimizes a unique linear filter for each subject and channel. After filtering, the spectra for each channel across all subjects are aligned to a common spectrum, which is the barycenter of the training subjects’ spectra for that channel. Originally, CMMN was applied to sleep staging with deep neural networks. In this work, we extend it to the case of automatically labeling independent components, such as the popular ICLabel classifier [14], which can support automatic EEG artifact removal and denoising. In comparison to the ICLabel classifier, which uses a neural network model with scalp map and spectral features (including the autocorrelation sequence), and was trained on an extensive dataset [15], we consider custom classifiers trained on a much smaller dataset collected in the US [13] using only time-series features extracted from each independent component [11]—that is, without any spatial information. We then test on an another independent dataset [8], collected in Europe. These two datasets were used in a prior work on IC classification [5], and are now both available on OpenNeuro [12, 9]. Because these datasets have expert-labeled ICs, we avoid filtering each channel with a different filter, which would change each IC differently; instead, we find a single CMMN filter for each subject that is applied to all channels. This channel-averaged CMMN filter is based on matching the channel-averaged spectrum of each test subject to either the barycenter of the channel-averaged spectra from the training set or the nearest spectrum in the training set in terms of the optimal transport distance. With a shared single temporal filter across channels, the filtering can be applied before or after the spatial filtering achieved by ICA’s demixing matrix. Crucially, the use of the channel-averaging enables the application of CMMN to target domains with a differing number of channels. Here, the source (train) and target (test) datasets have 134–235 channels and 64 channels, respectively. This extension of the CMMN methodology achieves domain adaptation between two datasets with the pre-trained classifiers such that their performance (the subject averaged F1 score for the brain class is 0.91) exceeds the performance of ICLabel (0.88).
1 Methodology
@@
The CMMN methodology enables domain adaptation between different neural signals [7] by mapping the spectra of the target subjects to the barycenter of the source subjects via a linear filter that achieves the optimal transport with respect to an $\ell_{2}$-metric, assuming the signals are zero-mean stationary Gaussian discrete-time processes (see Appendix A for more details). We extend this methodology in two key ways: we propose to use channel-averaged PSDs yielding a single filter per subject and a subject-to-subject mapping scheme. A key benefit of using a single filter per subject is that it enables adapting between EEG datasets with a variable number of channels. We also propose to use $\ell_{1}$-normalization of spectra for invariance to signal scale, which may occur due to impedance differences due to electrodes or scalp contact or differences in the electronic systems (amplifiers and filtering) or digital filtering.
@@
We assume access to data (multichannel EEG) from $I$ subjects in the source domain and one or more target subjects. We apply CMMN to transform the target EEGs to the source domain and apply classifiers trained on the source domain to the transformed target data. Let $x_{c}[n]$ denote the $c$-th channel of the target signal, $c\in\{1,\ldots,C^{\text{T}}\}$. The CMMN transformation is a linear filtering of $x_{c}[n]$, with output $y_{c}[n]=(h*x_{c})[n],\quad c\in\{1,\ldots,C^{\text{T}}\}$, where $h[n]$ is the normalizing CMMN filter, and $y_{c}[n]$ is the transformed signal that will have a source-like spectrum (the channel-averaged PSDs will match). The estimation of the normalizing CMMN filter is a three-step process that uses a reference spectrum (computed in the second step) that is either a Barycenter from the source or the normalized PSD of a source subject via the subject-to-subject (Subj-to-subj) assignment: 1) power spectral density calculation, 2a) barycenter calculation, 2b) minimization of the Hellinger distance between target signals and source signals for the Subj-to-subj mapping scheme, and 3) calculation of the normalizing filters.
@@
Step 1 PSD calculation. As in [7], we use the Welch periodogram method [17], which takes averages of the squared FFT from possibly overlapping windows, to calculate PSDs for each subject, as implemented by the SciPy library [16]. For real-valued signals, the spectra will be conjugate symmetric for positive and negative frequencies. In this case, $\bm{p}\in\mathbb{R}_{\geq 0}^{P}$ denotes the PSD for $P$ non-negative frequencies (for zero-mean signals the DC component should be zero but is retained). For $K$ equal-sized windows of length $M$, $\{\bm{x}_{k}\}_{k=1}^{K}=\mathcal{X}\subset\mathbb{R}^{M}$ with windowing function $\bm{w}$, the power spectral density estimate is $\bm{p}=\frac{1}{|\mathcal{X}|}\sum_{\bm{x}\in\mathcal{X}}\lvert\mathrm{RFFT}(\bm{w}\odot\bm{x})\rvert^{\odot 2}\in\mathbb{R}_{\geq 0}^{P}$, where $\mathrm{RFFT}$ denotes the FFT operation for real-valued signals and returns the density estimate only for non-negative frequencies, $P=\mathtt{nfft}/2+1$ is the number of non-negative frequencies, and $N=\mathtt{nfft}$ is the length of the windows after zero-padding, assumed to be even. For a subject with $C$ channels, the matrix of PSDs is $\bm{P}=[\bm{p}_{1},\ldots,\bm{p}_{C}]^{\top}\in\mathbb{R}^{C\times P}$. Unlike the original CMMN, which used different filters for each channel, we propose to compute a channel-averaged PSD as $\bar{\bm{p}}=\frac{1}{C_{i}}\sum_{c=1}^{C}\bm{p}_{c}$. [^1] One possible concern is whether channel-averaging is a sufficient characterization of multi-channel EEG given the spectral differences in signals across the scalp. Of course, if the location of the electrodes on the scalp is vastly different such that the spectral content is not comparable, then channel-averaged PSD may not be meaningful, but for EEG montages with whole scalp this should not be an issue.
@@
Step 2a Barycenter calculation. We compute a barycenter for the source subjects only, this serves as the template for mapping the target subjects. For $I$ source subjects with channel-averaged PSDs $\bar{\bm{p}}_{1}^{\text{S}},\ldots,\bar{\bm{p}}_{I}^{\text{S}}$, the barycenter, $\bar{\bm{p}}_{\text{S}}$, and the $\ell_{1}$-normalized barycenter, $\tilde{\bm{p}}_{\text{S}}$, are defined as
${\bar{\bm{p}}}_{\text{S}}=\frac{1}{I}\sum_{i=1}^{I}\bar{\bm{p}}^{\text{S}}_{i},\quad{\tilde{\bm{p}}}_{\text{S}}=\frac{1}{I}\sum_{i=1}^{I}\frac{\bar{\bm{p}}^{\text{S}}_{i}}{\|\bar{\bm{p}}^{\text{S}}_{i}\|_{1}}=\frac{1}{I}\sum_{i=1}^{I}\tilde{\bm{p}}^{\text{S}}_{i},\quad\tilde{\bm{p}}^{\text{S}}_{i}=\frac{\bar{\bm{p}}^{\text{S}}_{i}}{\|\bar{\bm{p}}^{\text{S}}_{i}\|_{1}},$ (1)
@@
The $l_{1}$ normalization ensures that each subject contributes equally to the average. As the units of PSD are squared, outliers without normalization can influence the average.
Step 2b: Subject-to-Subject Mapping. As an alternative to the barycenter scheme, we map a target subject to the nearest source subject. For a target subject with an $\ell_{1}$-normalized PSD $\tilde{\bm{p}}^{\text{T}}=\frac{\bar{\bm{p}}^{\text{T}}}{\|\bar{\bm{p}}^{\text{T}}\|_{1}}$, we find the index $i^{*}$ of the source subject that minimizes the Hellinger distance, which is applicable since the $\ell_{1}$-normalized PSD is a probability mass function (PMF)—non-negative and sums to 1. The channel-averaged PSD of the best-matched source is then $\bar{\bm{p}}_{*}^{\text{S}}=\bar{\bm{p}}_{i^{*}}^{\text{S}}$, where
@@
$i^{*}=\operatorname*{arg\,min}_{i\in\{1,\ldots,I\}}d_{\text{He}}(\tilde{\bm{p}}^{\text{S}}_{i},\tilde{\bm{p}}^{\text{T}})=\frac{1}{\sqrt{2}}\left\lVert(\tilde{\bm{p}}^{\text{S}}_{i})^{\odot\frac{1}{2}}-(\tilde{\bm{p}}^{\text{T}})^{\odot\frac{1}{2}}\right\rVert_{2}=\sqrt{\frac{1}{2}\sum_{n=0}^{N-1}\left(\sqrt{\tilde{p}_{i}^{\text{S}}[n]}-\sqrt{\tilde{p}^{\text{T}}[n]}\right)^{2}},$ (2)
where $d_{\text{He}}$ is the Hellinger distance. Assuming that $\tilde{\bm{p}}^{\text{S}}_{i}$, $\tilde{\bm{p}}^{\text{T}}$ are the eigenvalues of the trace normalized circulant matrices $\bm{\rho}^{\text{S}}_{i}=\frac{\bm{\Sigma}^{\text{S}}_{i}}{\mathrm{tr}(\bm{\Sigma}^{\text{S}}_{i})}$ and $\bm{\rho}^{\text{T}}=\frac{\bm{\Sigma}^{\text{T}}}{\mathrm{tr}(\bm{\Sigma}^{\text{T}})}$, where $\bm{\Sigma}^{\text{S}}_{i}$ and $\bm{\Sigma}^{\text{T}}$ are the covariance matrices of the discrete-time zero-mean stationary Gaussian random processes, then the Wasserstein-2 distance between the two variance-normalized processes is
$W_{2}($ $\mathcal{N}(\bm{0},\bm{\rho}^{\text{S}}_{i}),\mathcal{N}(\bm{0},\bm{\rho}^{\text{T}}))=\sqrt{2}d_{\text{He}}(\tilde{\bm{p}}^{\text{S}}_{i},\tilde{\bm{p}}^{\text{T}}).$ (3)
The use of Hellinger distance compared to Bures-Wasserstein distance means that signals are compared in terms of their spectral shape, ignoring difference in their variance.
@@
Step 3 Normalizing filter. We define a real-valued, zero-phase linear filter with impulse response $h[n]$ and frequency response $\bm{H}$ given by the square root of the ratio of the channel-averaged PSDs [4]:
@@
$\bm{h}$ $=\mathrm{IRFFT}_{M}(\bm{H}),\quad\bm{H}=\bar{\bm{p}}^{\text{T}\odot-\frac{1}{2}}\odot\bar{\bm{p}}^{\text{S}\odot\frac{1}{2}},\quad H[n]=\frac{\sqrt{\bar{p}^{\text{S}}[n]}}{\sqrt{\bar{p}^{\text{T}}[n]}},n\in\{0,\ldots,P-1\},$ (4)